$SPAD/src/input f04atf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)cl all
showArrayValues true
showScalarValues true
a:Matrix SF:=
 [[33,16,72],_
  [-24,-10,-57],_
  [-8,-4,-17],_
  [0,0,0],_
  [0,0,0],_
  [0,0,0],_
  [0,0,0],_
  [0,0,0]]
ia:=8
b:Matrix SF:=
 [[-359 ,281 ,85 ]]
n:=3
iaa:=8
result:=f04atf(a,ia,b,n,iaa,-1)

———-

References^{[1] nothing}

4